# Optimal convergence rates in trace distance and relative entropy for the quantum central limit theorem

### v1

@@ 3.1 Moments of a quantum state
Let $\rho$ be a quantum state and $\kappa>0$. Following [3], we say that $\rho$ has finite moments up to order $\kappa$ if
@@
where $\|z\|^{2}=\sum_{i=1}^{m}|z_{i}|^{2}$. As a result, for any quantum state $\rho$, we have
@@
Therefore, it is sufficient to prove the lemma for $m=1$. In fact, we can set $g_{m,\kappa}(\cdot)=m^{\kappa/2-1}g_{1,\kappa}(\cdot)$. Furthermore, for $m=1$, it suffices to establish the theorem for even values of $\kappa$ since, assuming the existence of a $g_{1,\kappa+1}(\|z\|)$ satisfying the lemma for $\kappa+1$, we have for $10<\lambda=\frac{\kappa}{\kappa+1}<$ 1 that
@@
By (20), the characteristic function of $\rho$ is expressed in terms of the characteristic function of $\sigma$. Then, using classical results that provide bounds on derivatives of the characteristic function in terms of absolute moments, it is shown in [3] that for any positive integer $\kappa$ and $\epsilon>0$, there exists a constant $c^{\prime\prime}(\kappa,\epsilon)$ such that for any $m$-mode quantum state $\rho$, and any $\alpha,\beta\in\mathbb{R}_{+}^{m}$ with $|\alpha|+|\beta|\leq\kappa$, we have
@@ 3.2 Taylor expansion of the characteristic function
Let $\rho$ be a centered $m$-mode quantum state with finite $\kappa$-th-order moment where $\kappa\geq 4$. By equation (21), we have
@@
$|\alpha|=\sum_{i=1}^{2m}\alpha_{i},\quad z^{\alpha}=\prod_{i=1}^{m}z_{i}^{\alpha_{2i-1}}\bar{z_{i}}^{\alpha_{2i}},\quad\alpha!=\alpha_{1}!\dots\alpha_{2m}!.$ (34)
@@
However, in order to prove stronger versions of the quantum CLT, we need to carefully analyze the expansion (36). In the rest of this section, we will present two important lemmas that will aid in proving the optimal convergence rates for the quantum CLT under minimal assumptions.
@@
Let $\rho$ be an $m$-mode quantum state that is centered and has a finite $\kappa$-th-order moment for some integer $\kappa\geq 4$, and let $\nu_{\min}=\|\boldsymbol{\gamma}(\rho)^{-1}\|^{-1}$ and $\nu_{\max}=\|\boldsymbol{\gamma}(\rho)\|$ be the minimum and maximum eigenvalues of the covariance matrix of $\rho$, respectively. Then, there exist a constant $C=C(\nu_{\min},\nu_{\max})>0$ and a polynomial $L(\cdot)=L_{\nu_{\min},\nu_{\max}}(\cdot)$ whose degree depends only on $\kappa,\alpha,\beta$, and whose coefficients depend only on $\nu_{\min},\nu_{\max}$, such that the following holds: for any $z\in\mathbb{C}^{m}$ with $\|z\|\leq CM_{\kappa}(\rho)^{-1/(\kappa-2)}\sqrt{n}$, and any $\alpha,\beta\in\mathbb{Z}_{+}^{m}$ with $|\alpha|+|\beta|\leq\kappa$, we have
@@ 4 State truncation via Fock basis
To prove Theorem 1 and Theorem 2 under the assumption of finite moments of constant order independent of $m$, the number of modes, we need to approximate the underlying quantum state $\rho$ with states that also possess finite higher-order moments. To achieve this, we introduce a method we refer to as state truncation via the Fock basis, which constructs a sequence of finite-rank states $\sigma_{n}$ that approximates $\rho$, ensuring that all moments of $\sigma_{n}$ are finite. However, since our truncated states $\sigma_{n}$ depend on $n$, their moments will also vary with $n$. Therefore, our analysis requires careful consideration of the moments of $\sigma_{n}$ as well as its characteristic function.
@@
Note that all moments of $\sigma_{n}$ are finite; however, they depend on $n$. The following lemma helps ensure that the trace distance between the $n$-fold convolution of $\rho$ and the $n$-fold convolution of $\sigma_{n}$ decays at a sufficient rate, allowing us to use them to prove Theorem 1 and Theorem 2.
@@
Let $\rho$ be an $m$-mode quantum state with finite $s$-th order moment for some $s\geq 3$. With the above notation, we have
@@
where in the last line we use the fact that the $s$-th order moment of $\rho$ is finite. Using this we obtain
@@
It is important to note that the states $\sigma_{n}$ are generally not centered. In our arguments, we apply displacement operators to center them, and as we will see, it is beneficial to have upper bounds on the first moments of $\sigma_{n}$. The following lemma provides us with such an upper bound.
@@
Let $\rho$ be a centered $m$-mode quantum state with finite $s$-th order moment for some $s\geq 3$. Then, for any $j$ we have $|z_{n,j}|=\mathcal{O}(\frac{1}{n^{(s-1)/2}})$ where $z_{n,j}=\mathrm{tr}(\sigma_{n}\mathbf{a}_{j})$.
@@
In the first line, we utilize the fact that, according to (41), for large $n$, we have $\mathrm{tr}(\rho_{n})\geq\frac{1}{2}$. The second line leverages the fact that $\rho$ is centered. The fourth line follows from the Cauchy–Schwarz and AM-GM inequalities, while in the sixth line, we use the fact that $|k|\geq k_{j}$. Finally, the last line follows from the finiteness of the $s$-th order moment of $\rho$. ∎
@@
Let $\rho$ be a centered $m$-mode quantum state with finite $s$-th order moment for some $s\geq 3$. With the above notation, we have
@@
For the first term, by (41), we have $\frac{1-\mathrm{tr}(\rho_{n})}{\mathrm{tr}(\rho_{n})}=\mathcal{O}(n^{-s/2})$. Applying the Cauchy–Schwarz inequality yields $\sum_{|k|\leq n}\sqrt{k_{j}k_{j^{\prime}}}\big{|}\langle k|\rho|k^{-j,-{j^{\prime}}}\rangle\big{|}<+\infty$ since the second-order moment of $\rho$ is finite. For the second term, by applying a similar chain of inequalities as in the proof of Lemma 8 and using the finiteness of the $s$-th order moment, we obtain
@@
Combining these, we establish the desired result. ∎
@@
Let $\rho$ be a centered $m$-mode quantum state with finite $s$-th order moment for some $s\geq 3$. For any $\kappa\geq s$, we have $M_{\kappa}(\tilde{\sigma}_{n})=\mathcal{O}\big{(}M_{s}(\rho)n^{(\kappa-s)/2}\big{)}$.
@@
Let $\rho$ be a centered $m$-mode state with finite $s$-th order moment such that $s\geq 3$. For any $\epsilon>0$, there are constants $\theta_{0},\theta_{1}>0$ such that for any sufficiently large $n$ and any $\|z\|\leq\epsilon\sqrt{n}$ we have
@@
We note that there is constant $b>0$ such that $\hat{z}^{\dagger}\Lambda_{m}^{\dagger}\boldsymbol{\gamma}(\tilde{\sigma}_{n})\Lambda_{m}\hat{z}\leq bM_{2}(\tilde{\sigma}_{m})\|z\|^{2}\leq bM_{3}(\tilde{\sigma}_{m})^{2/3}\|z\|^{2}$ where the second inequality follows from Lemma 2. Therefore,
@@
$\leq\exp\Big{(}2m\theta_{0}-\frac{1}{4}\hat{z}^{\dagger}\Lambda_{m}^{\dagger}\boldsymbol{\gamma}(\tilde{\sigma}_{n})\Lambda_{m}\hat{z}+\theta^{\prime}_{1}M_{3}(\tilde{\sigma}_{n})\epsilon\|z\|^{2}\Big{)}$
@@ 5 An upper bound on the quantum relative entropy
As an immediate consequence of this result, we derive an upper bound for the relative entropy of non-Gaussianity of a state $\rho$, symbolized as $d_{\mathcal{G}}(\rho)$. This quantity, defined as the relative entropy distance between $\rho$ and the collection $\mathcal{G}$ of all Gaussian states [18, 32], serves as a metric for non-Gaussian features of quantum states:
@@
The measure $d_{\mathcal{G}}(\rho)$ has several advantageous properties, making it a valuable indicator of non-Gaussianity. For instance, it is faithful: $d_{\mathcal{G}}(\rho)\geq 0$, reaching zero exclusively when $\rho$ is a Gaussian state. The minimum value in this definition is achieved by the Gaussian reference state $\rho_{G}$, so that
@@
To establish Theorem 3, we start by proving two auxiliary lemmas that will be used in its proof. Initially, we assume that $\tau$ is a thermal state, as specified in the theorem statement. Subsequently, we will address the proof in the general case.
@@
Since $E(z)$ is an odd polynomial, $e_{p,q}\neq 0$ only if $\sum_{j}(p_{j}+q_{j})$ is an odd number not exceeding the total degree of $E(z)$. Note that to write down the above expression we in particular use $\mathbf{a}_{j}\mathbf{a}_{j}^{\dagger}=\mathbf{a}_{j}^{\dagger}\mathbf{a}_{j}+1$, yet this would not increase the degrees of the terms on the right hand side nor it changes their parity. This representation of $\tau_{\alpha}$ in particular implies that the diagonal entries of $\tau_{\alpha}$ with respect to the Fock basis are independent of $\alpha$:
@@
$=\frac{(\beta_{\max}+1)e^{t}}{\nu_{\beta}(t+1)^{m}}\mathrm{tr}\big{(}(\rho-\tau_{\alpha})^{2}(N_{m}+m)^{m}\big{)}$
To estimate $\sum_{\beta\cdot k\leq t}\langle k|\tau^{-\frac{1}{2}}({\tau_{\alpha}}-\tau)^{2}\tau^{-\frac{1}{2}}|k\rangle$, we let $\Delta:=\frac{1}{\alpha}({\tau_{\alpha}}-\tau)$. Invoking (52) we have
@@
where $M_{\beta,E}>0$ is a constant depending only on $\beta$ and the polynomial $E(z)$. Note that $M_{\beta,E}$ is finite since $\tau$ is Gaussian and all of whose moments are finite.
We conclude that
@@
$C^{\prime\prime}_{\beta}=\frac{2^{m}m\eta_{\beta}(\beta_{1}+1)^{2}\cdots(\beta_{m}+1)^{2}}{\nu_{\beta}}.$
@@ 6 Proof of Theorem 1
Since the case $m=1$ is already proven in [4], we assume $m\geq 2$. By applying a suitable Gaussian unitary and using (19), we may, without loss of generality, assume that $\rho$ is in Williamson form with a diagonal covariance matrix as in (16). Let $\nu_{\min}$ and $\nu_{\max}$ denote the minimum and maximum eigenvalues of $\boldsymbol{\gamma}(\rho)$, respectively. As discussed in Subsection 2.2, if $\nu_{j}=1$ for some $j$, then the $j$-th mode of $\rho$ is in the vacuum state and decouples from the other modes. Since the $n$-fold convolution of the vacuum remains the vacuum, this mode does not contribute to the trace distance $\big{\|}\rho^{\boxplus n}-\rho_{G}\big{\|}_{1}$. Consequently, we can disregard the $j$-th mode in the proof of Theorem 1. Thus, we assume in the following that the minimum eigenvalue of the covariance matrix is strictly greater than one:
@@
Applying the triangle inequality, by Lemma 7 we have
@@
Invoking Lemma 1 with $T_{n}:=\rho_{G}-\tilde{\sigma}_{n,G}$, we get
@@
Since $\rho$ is in Williamson form, the Gaussian state $\rho_{G}$ is thermal and is equal to $\rho_{G}=\tau_{1}\otimes\cdots\otimes\tau_{n}$ where
@@
where $\beta_{j}>0$ is given by
@@
To prove this inequality we once again use Lemma 1. Note that all moments of $\tilde{\sigma}_{n}$ and $\tilde{\sigma}_{n,G}$ are finite, yet they depend on $n$. Define $T_{n}=\tilde{\sigma}_{n}^{\boxplus n}-\tilde{\sigma}_{n,G}$. Let
@@
First term: To handle the first term we use (60).
To bound the derivative on the right hand side of (60), we apply Lemma 5. Using this lemma for $\kappa=2m$, for sufficiently large $n$ and any $z$ with $\|z\|\leq\epsilon_{n}\sqrt{n}$, we have
@@
where the sum runs over all tuples $k=(k_{1},\dots,k_{m})$ and $\ell=(\ell_{1},\dots,\ell_{m})$ satisfying $1\leq k_{j},\ell_{j^{\prime}}\leq 2m$, and the coefficients $f_{k,\ell}$ are at most $|f_{k,\ell}|=\mathcal{O}(n^{2m})$. By (23), there exists $\delta_{n}>0$ such that $\big{|}\chi_{\sigma_{n}}\big{(}z/\sqrt{n}\big{)}\big{|}\leq 1-\delta_{n}$ for all $\|z\|\geq\epsilon\sqrt{n}$. Note that $\delta_{n}$ is determined in terms of the second moment of $\tilde{\sigma}_{n}$ (see  [3, Proposition 15] for the exact form of $\delta_{n}$), and by Lemma 9 for sufficiently large $n$ the second moment of $\tilde{\sigma}_{n}$ can be bounded in terms of the second moment of $\rho$. Hence, there exists a $\delta>0$ such that for sufficiently large $n$, we have $\delta_{n}\geq\delta$. Thus, for sufficiently large $n$, we have
@@
Note that by definition $M_{4m}\big{(}\tilde{\sigma}_{n}^{\otimes 2m}\big{)}=\mathcal{O}\big{(}M_{4m}(\tilde{\sigma}_{n})\big{)}$, and by (69), we have $M_{4m}(\tilde{\sigma}_{n})=\mathcal{O}(n^{(4m-3)/2})$. Substituting these bounds into (74), we find that the last term in (66) is exponentially small.

### v2

@@ 3.1 Moments of a quantum state
Let $\rho$ be a quantum state and $\kappa>0$. Following [3], we say that $\rho$ has finite moments up to order $\kappa$ if [^1] The moment of order $\kappa$ is sometimes defined as $\mathrm{tr}\Big{(}\rho N_{m}^{\frac{\kappa}{2}}\Big{)}$ in the literature. We note, however, that the finiteness of moments of order $\kappa$ according to these two definitions coincide as we have $\mathrm{tr}\Big{(}\rho N_{m}^{\frac{\kappa}{2}}\Big{)}\leq M_{\kappa}(\rho)\leq 2^{\frac{\kappa}{2}}\mathrm{tr}\Big{(}\rho N_{m}^{\frac{\kappa}{2}}\Big{)}+1$
@@
where $\|z\|^{2}=\sum_{i=1}^{m}|z_{i}|^{2}$. As a result, for any $m$-mode quantum state $\rho$, we have
@@
Therefore, it is sufficient to prove the lemma for $m=1$. In fact, we can set $g_{m,\kappa}(\cdot)=m^{\kappa/2-1}g_{1,\kappa}(\cdot)$. Furthermore, for $m=1$, it suffices to establish the theorem for even values of $\kappa$ since, assuming the existence of a polynomial $g_{1,\kappa+1}(\|z\|)$ satisfying the lemma for $\kappa+1$, we have for $10<\lambda=\frac{\kappa}{\kappa+1}<$ 1 that
@@
By (21), the characteristic function of $\rho$ is expressed in terms of the characteristic function of $\sigma$. Then, using classical results that provide bounds on the derivatives of the characteristic function in terms of absolute moments, it is shown in [3] that for any positive integer $\kappa$ and $\epsilon>0$, there exists a constant $c^{\prime\prime}(\kappa,\epsilon)$ such that for any $m$-mode quantum state $\rho$, and any $\alpha,\beta\in\mathbb{R}_{+}^{m}$ with $|\alpha|+|\beta|\leq\kappa$, we have
@@ 3.2 Taylor expansion of the characteristic function
Let $\rho$ be a centered $m$-mode quantum state with finite $\kappa$-th-order moments where $\kappa\geq 4$. By equation (22), we have
@@
$|\alpha|=\sum_{i=1}^{2m}\alpha_{i},\quad z^{\alpha}=\prod_{i=1}^{m}z_{i}^{\alpha_{2i-1}}\bar{z_{i}}^{\alpha_{2i}},\quad\alpha!=\alpha_{1}!\cdots\alpha_{2m}!.$ (35)
@@
However, in order to prove stronger versions of the quantum CLT, we need to carefully analyze the quantum Edgeworth-type expansion (37). In the rest of this section, we will present two important lemmas that will aid in proving the optimal convergence rates for the quantum CLT under minimal assumptions.
@@
Let $\rho$ be an $m$-mode quantum state that is centered and has finite $\kappa$-th-order moments for some integer $\kappa\geq 4$, and let $\nu_{\min}=\|\boldsymbol{\gamma}(\rho)^{-1}\|^{-1}$ and $\nu_{\max}=\|\boldsymbol{\gamma}(\rho)\|$ be the minimum and maximum eigenvalues of the covariance matrix of $\rho$, respectively. Then, there exist a constant $C=C(\nu_{\min},\nu_{\max})>0$ and a polynomial $L(\cdot)=L_{\nu_{\min},\nu_{\max}}(\cdot)$ whose degree depends only on $\kappa,\alpha,\beta$, and whose coefficients depend only on $\nu_{\min},\nu_{\max}$, such that the following holds: for any $z\in\mathbb{C}^{m}$ with $\|z\|\leq CM_{\kappa}(\rho)^{-1/(\kappa-2)}\sqrt{n}$, and any $\alpha,\beta\in\mathbb{Z}_{+}^{m}$ with $|\alpha|+|\beta|\leq\kappa$, we have
@@ 4 State truncation via Fock basis
To prove Theorem 1 and Theorem 2 under the assumption of finite moments of constant order independent of $m$, the number of modes, we need to approximate the underlying quantum state $\rho$ with states that also possess finite higher-order moments. To achieve this, we introduce a method we refer to as state truncation via the Fock basis, which constructs a sequence of finite-rank states $\sigma_{n}$ that approximate $\rho$, ensuring that all the moments of $\sigma_{n}$ are finite. However, since our truncated states $\sigma_{n}$ depend on $n$, their moments will also vary with $n$. Therefore, our analysis requires careful consideration of the moments of $\sigma_{n}$ as well as its characteristic function.
@@
Note that all moments of $\sigma_{n}$ are finite, yet they depend on $n$. The following lemma helps to ensure that the trace distance between the $n$-fold convolution of $\rho$ and the $n$-fold convolution of $\sigma_{n}$ decays at a sufficient rate, allowing us to use them to prove Theorem 1 and Theorem 2.
@@
Let $\rho$ be an $m$-mode quantum state with finite $s$-th order moments for some $s\geq 3$. With the above notation, we have
@@
where in the last line we use the fact that the $s$-th order moments of $\rho$ are finite. Using this we obtain
@@
It is important to note that the states $\sigma_{n}$ are generally not centered. In our arguments, we apply displacement operators to center them, so as we will see, it is beneficial to have upper bounds on the first moments of $\sigma_{n}$. The following lemma provides us with such an upper bound.
@@
Let $\rho$ be a centered $m$-mode quantum state with finite $s$-th order moments for some $s\geq 3$. Then, for any $j$ we have $|z_{n,j}|=\mathcal{O}\big{(}\frac{1}{n^{(s-1)/2}}\big{)}$ where $z_{n,j}=\mathrm{tr}(\sigma_{n}\mathbf{a}_{j})$.
@@
In the first line, we utilize the fact that, according to (42), for large $n$, we have $\mathrm{tr}(\rho_{n})\geq\frac{1}{2}$. The second line leverages the fact that $\rho$ is centered. The fourth line follows from the Cauchy–Schwarz and AM-GM inequalities, while in the sixth line, we use the fact that $|k|\geq k_{j}$. Finally, the last line follows from the finiteness of the $s$-th order moments of $\rho$. ∎
@@
Let $\rho$ be a centered $m$-mode quantum state with finite $s$-th order moments for some $s\geq 3$. With the above notation, we have
@@
For the first term, by (42), we have $\frac{1-\mathrm{tr}(\rho_{n})}{\mathrm{tr}(\rho_{n})}=\mathcal{O}(n^{-s/2})$. Applying the Cauchy–Schwarz inequality yields $\sum_{|k|\leq n}\sqrt{k_{j}k_{j^{\prime}}}\big{|}\langle k|\rho|k^{-j,-{j^{\prime}}}\rangle\big{|}<+\infty$ since the second-order moments of $\rho$ are finite. For the second term, by applying a similar chain of inequalities as in the proof of Lemma 8 and using the finiteness of the $s$-th order moments, we obtain
@@
Combining these, we establish the desired bound. ∎
@@
Let $\rho$ be a centered $m$-mode quantum state with finite $s$-th order moments for some $s\geq 3$. For any $\kappa\geq s$, we have $M_{\kappa}(\tilde{\sigma}_{n})=\mathcal{O}\big{(}M_{s}(\rho)n^{(\kappa-s)/2}\big{)}$.
@@
Let $\rho$ be a centered $m$-mode state with finite $s$-th order moments such that $s\geq 3$. For any $\epsilon>0$, there are constants $\theta_{0},\theta_{1}>0$ such that for any sufficiently large $n$ and any $\|z\|\leq\epsilon\sqrt{n}$ we have
@@
We note that there is a constant $b>0$ such that
$\hat{z}^{\dagger}\Lambda_{m}^{\dagger}\boldsymbol{\gamma}(\tilde{\sigma}_{n})\Lambda_{m}\hat{z}\leq bM_{2}(\tilde{\sigma}_{m})\|z\|^{2}\leq bM_{3}(\tilde{\sigma}_{m})^{2/3}\|z\|^{2},$
where the second inequality follows from Lemma 2. Therefore,
@@
$\leq\exp\Big{(}2m\theta_{0}-\frac{1}{4}\hat{z}^{\dagger}\Lambda_{m}^{\dagger}\boldsymbol{\gamma}(\tilde{\sigma}_{n})\Lambda_{m}\hat{z}+\theta^{\prime}_{1}M_{3}(\tilde{\sigma}_{n})\epsilon\|z\|^{2}\Big{)}.$
@@ 5 An upper bound on the quantum relative entropy
As an immediate consequence of this result, we derive an upper bound for the relative entropy of non-Gaussianity of a state $\rho$, symbolized as $d_{\mathcal{G}}(\rho)$ [17, 32]. This quantity, defined as the relative entropy between $\rho$ and the collection $\mathcal{G}$ of all Gaussian states, serves as a metric for non-Gaussian features of quantum states:
@@
The measure $d_{\mathcal{G}}(\rho)$ has several advantageous properties, making it a valuable indicator of non-Gaussianity. For instance, it is faithful: $d_{\mathcal{G}}(\rho)\geq 0$, reaching zero exclusively when $\rho$ is a Gaussian state. In fact, the minimum value in this definition is achieved by the Gaussian reference state $\rho_{G}$, so that
@@
To establish Theorem 4, we start by proving two auxiliary lemmas. Initially, we assume that $\tau$ is a thermal state, as specified in the theorem statement. Subsequently, we will address the proof in the general case.
@@
Since $E(z)$ is an odd polynomial, $e_{p,q}\neq 0$ only if $\sum_{j}(p_{j}+q_{j})$ is an odd number not exceeding the total degree of $E(z)$. Note that to write down the above expression we in particular use $\mathbf{a}_{j}\mathbf{a}_{j}^{\dagger}=\mathbf{a}_{j}^{\dagger}\mathbf{a}_{j}+1$, yet this would not increase the degrees of the terms on the right hand side nor it changes their parity. This representation of $\tau_{\alpha}$ implies that the diagonal entries of $\tau_{\alpha}$ with respect to the Fock basis are independent of $\alpha$:
@@
$=\frac{(\beta_{\max}+1)e^{t}}{\nu_{\beta}(t+1)^{m}}\mathrm{tr}\big{(}(\rho-\tau_{\alpha})^{2}(N_{m}+m)^{m}\big{)}.$
We continue by letting $\sum_{\beta\cdot k\leq t}\langle k|\tau^{-\frac{1}{2}}({\tau_{\alpha}}-\tau)^{2}\tau^{-\frac{1}{2}}|k\rangle$, we let $\Delta:=\frac{1}{\alpha}({\tau_{\alpha}}-\tau)$. Then, invoking (53) we have
@@
where $M_{\beta,E}>0$ is a constant depending only on $\beta$ and the polynomial $E(z)$. Note that $M_{\beta,E}$ is finite since $\tau$ is Gaussian and all of whose moments are finite. We conclude that
@@
$C^{\prime\prime}_{\beta}=\frac{2^{m}m\eta_{\beta}(\beta_{1}+1)^{2}\cdots(\beta_{m}+1)^{2}}{\nu_{\beta}},$
@@ 6 Proof of Theorem 1
Since the case $m=1$ is already proven in [4], we assume $m\geq 2$. By applying a suitable Gaussian unitary and using (20), we may, without loss of generality, assume that $\rho$ is in the Williamson form with a diagonal covariance matrix as in (17). Let $\nu_{\min}$ and $\nu_{\max}$ denote the minimum and maximum eigenvalues of $\boldsymbol{\gamma}(\rho)$, respectively. As discussed in Subsection 2.2, if $\nu_{j}=1$ for some $j$, then the $j$-th mode of $\rho$ is in the vacuum state and decoupled from the other modes. Since the $n$-fold convolution of the vacuum remains the vacuum, this mode does not contribute to the trace distance $\big{\|}\rho^{\boxplus n}-\rho_{G}\big{\|}_{1}$. Consequently, we can disregard the $j$-th mode in the proof of Theorem 1. Thus, we assume in the following that the minimum eigenvalue of the covariance matrix is strictly greater than one:
@@
Using the notation developed in Section 4 and applying the triangle inequality, by Lemma 7 we have
@@
To prove this claim, invoking Lemma 1 with $T_{n}:=\rho_{G}-\tilde{\sigma}_{n,G}$, we get
@@
Since $\rho$ is in the Williamson form, the Gaussian state $\rho_{G}$ is thermal and is equal to $\rho_{G}=\tau_{1}\otimes\cdots\otimes\tau_{n}$ where
@@
with $\beta_{j}>0$ being given by
@@
To prove this inequality we once again use Lemma 1. Note that all the moments of $\tilde{\sigma}_{n}$ and $\tilde{\sigma}_{n,G}$ are finite, yet they depend on $n$. Define $T_{n}=\tilde{\sigma}_{n}^{\boxplus n}-\tilde{\sigma}_{n,G}$. Let
@@
First term: To handle the first term we use (61). To bound the derivative on the right hand side of (61), we apply Lemma 5. Using this lemma for $\kappa=2m$, for sufficiently large $n$ and any $z$ with $\|z\|\leq\epsilon_{n}\sqrt{n}$, we have
@@
where the sum runs over all tuples $k=(k_{1},\dots,k_{m})$ and $\ell=(\ell_{1},\dots,\ell_{m})$ satisfying $1\leq k_{j},\ell_{j^{\prime}}\leq 2m$, and the coefficients $f_{k,\ell}$ are at most $|f_{k,\ell}|=\mathcal{O}(n^{2m})$. By (24), there exists $\delta_{n}>0$ such that $\big{|}\chi_{\sigma_{n}}\big{(}z/\sqrt{n}\big{)}\big{|}\leq 1-\delta_{n}$ for all $\|z\|\geq\epsilon\sqrt{n}$. Note that $\delta_{n}$ is determined in terms of the second moments of $\tilde{\sigma}_{n}$ (see  [3, Proposition 15] for the exact form of $\delta_{n}$), and by Lemma 9 for sufficiently large $n$ the second moments of $\tilde{\sigma}_{n}$ can be bounded in terms of the second moments of $\rho$. Hence, there exists a $\delta>0$ such that for sufficiently large $n$, we have $\delta_{n}\geq\delta$. Thus, for sufficiently large $n$, we have
@@
Note that by definition $M_{4m}\big{(}\tilde{\sigma}_{n}^{\otimes 2m}\big{)}=\mathcal{O}\big{(}M_{4m}(\tilde{\sigma}_{n})\big{)}$, and by (70), we have $M_{4m}(\tilde{\sigma}_{n})=\mathcal{O}\big{(}n^{(4m-3)/2}\big{)}$. Substituting these bounds into (75), we find that the last term in (67) is exponentially small.